OpenStudy (anonymous):

Find the vertex in each equation. Show your work, or explain your process. g(x)=1/2(x+3)^2-5 h(x)=x^2+4x-3

OpenStudy (anonymous):

The vertex for g(x) is (-3,-5) and the vertex for h(x) is (-2, -7)

OpenStudy (anonymous):

@aodnoah706

OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

yassssss naenaehoowaahh <3

OpenStudy (anonymous):

how did you find this out? work?

OpenStudy (anonymous):

When an equation is in vertex form f(x)=y = a(x – h)2 + k then you can find the vertex. h=x and k=y

OpenStudy (anonymous):

So g(x)=1/2(x+3)^2-5 has a vertex of (-3,-5) because -5 is k ans -3 is h because when there is a plus sign that means the negative signs canceled out, which means it started out as -3.

OpenStudy (anonymous):

Do you follow @aodnoah706 ?

OpenStudy (anonymous):

I found g(x) by using the formula -b/2a to get an x intercept of -2. -4/2(1)=-2

OpenStudy (anonymous):

i got the y intercept like this and solved for y. x^2+4x-3=y (-2)^2+4(-2)-3=y 4-8-3=y -7=y

OpenStudy (anonymous):

Do you get this @aodnoah706 ?